OpenStudy (anonymous):

Are the following points on the graph of f(x)=3x^2 + 7x -8 (-3,-2) (-1, -10) (1,2) (3, 45)

OpenStudy (lucaz):

for the first point (-1,-10) if you replace x by -1 and get -10 the point is on the graph

OpenStudy (lucaz):

it's not the first point, but anyway..

OpenStudy (anonymous):

how can you test and see if the points are on the line

OpenStudy (lucaz):

f(-1) = 3(-1)^2+7(-1)-8 f(-1) = 3-7-8 f(-1) = -12 the point (-1,-10) is not on the graph

OpenStudy (lucaz):

let's say you have the point (a,b) you know the point is on the line if f(a)=b

OpenStudy (lucaz):

the values must satisfy the equation

OpenStudy (anonymous):

even if they are points...ex. (5, -2)

OpenStudy (lucaz):

any point, you substitute x and calculate y, if they match then the point is on the line

OpenStudy (lucaz):

try the point (-3,-2) what is f(-3)?

OpenStudy (anonymous):

if plugged into the equation i originally provided ( f(x) = 3x^2 +7x -8... then it would be -2

OpenStudy (anonymous):

ohhhh

OpenStudy (lucaz):

this means the point (-3,-2) is on the line of this function

OpenStudy (anonymous):

THANK YOU SO MUCH! you really helped me put that into perspective without completely giving me the answer! thank you!

OpenStudy (lucaz):